OpenStudy (anonymous):

Homogeneous equation: (a/b)y'(t)+y(t) = x(t) the answer is y(t) = Ae^(-b/a)t Huh?

OpenStudy (anonymous):

i know its a first order differential

OpenStudy (anonymous):

but where does the "A" and minus sign come from?

OpenStudy (jamesj):

The equation in the general form in which you've written it is inhomogeneous. The homogeneous version of that equation is x(t) = 0. I.e., (a/b)y'(t)+y(t) = 0 Now this equation is separable. Separate and solve.

OpenStudy (jamesj):

I.e., dy/dx = -b/a . y dy/y = -b/a dx => ln y = -(b/a)x + C You take it from here.

OpenStudy (anonymous):

jamesj... i dont get this step: dy/dx = -b/a . y

OpenStudy (jamesj):

(a/b) y' + y = 0 (a/b) y' = -y y' = -(b/a)y dy/dx = -(b/a)y

OpenStudy (jamesj):

Hence \[ \int \frac{dy}{y} = \int -(b/a) dx \ \ \ \implies \ln y = -(b/a)x + C \]

OpenStudy (anonymous):

oke... y = e^-(b/2) + C but the "A" where does that one come from

OpenStudy (jamesj):

\[ \implies y = \exp( -(b/a)x + C ) \] Hence y = ...

myininaya (myininaya):

\[\frac{a}{b}y'+y=t\] now I want the coefficient of y' to be 1 so I will multiply both sides by b/a \[y'+\frac{b}{a}y=t \frac{b}{a}\] now i will multiply both sides by v \[vy'+\frac{b}{a} v y=t v \frac{b}{a}\] I want to choose this v such that it satisifies the following equation \[v'=\frac{b}{a}v\] I want to choose it this way so that i can write \[vy'+v'y=t v \frac{b}{a}\] so then I can write \[(vy)'=t v \frac{b}{a}\] so what is this v well we need to solve \[v'=\frac{b}{a}v\] we can do this by separation of variables \[\frac{dv}{dt} =\frac{b}{a} v\] \[\frac{1}{v} dv =\frac{b}{a} dt\] now integrate both sides \[\ln|v|=\frac{b}{a} t+k \text{ let k=0 }\] let v>0 so we can just write \[\ln(v)=\frac{b}{a}t\] \[v=e^{\frac{b}{a}t}\] so we have \[(e^{ \frac{b}{a}t}y)'=t e^{ \frac{b}{a} t}\frac{b}{a}\] now integrating both sides gives \[e^{ t \frac{b}{a}} y+c_1=\int\limits_{}^{}\frac{b}{a} t e^{\frac{b}{a} t} dt\] so now we have to figure out how to do \[\frac{b}{a}\int\limits_{}^{}t e^{\frac{ b}{a} t} dt =\frac{b}{a} [ t \frac{a}{b}e^{\frac{b}{a} t} -\int\limits_{}^{}\frac{a}{b} e^{\frac{ b}{a} t} dt]\] \[=te^{\frac{b}{a} t} -\frac{a}{b}e^{\frac{b}{a} t} +c_2\] so we have \[e^{ t \frac{b}{a}} y+c_1=te^{\frac{b}{a} t} -\frac{a}{b}e^{\frac{b}{a} t} +c_2\] now we only need to really write one constant since constants are closed under addition so we have \[e^{t \frac{b}{a} } y=te^{ \frac{b}{a} t}-\frac{a}{b} e^{\frac{b}{a}t}+C\]

myininaya (myininaya):

did i make a mistake somewhere?

OpenStudy (jamesj):

Your equation y = e^-(b/2) + C is wrong. \[ y = e^{-(b/a)x + C} = e^C.e^{-(b/a)x} = Ae^{-(b/a)x} \] where A = e^C

myininaya (myininaya):

no wait i think it looks good :)

OpenStudy (jamesj):

@myin, you solved a different equation. The homogeneous equation is (a/b) y' + y = 0.

OpenStudy (anonymous):

owhh i forgot the (...)

myininaya (myininaya):

oh

myininaya (myininaya):

oops i wrote it down wrong altogether

OpenStudy (anonymous):

jamesj... the \[\int\limits_{}^{} -(b/a)dx\] isnt it \[-(b/a) (x + c)\]

myininaya (myininaya):

thats right

myininaya (myininaya):

you can also write it as -bx/a+c

OpenStudy (jamesj):

It can be, in which case you're just scaling the constant c. It makes no difference to the final answer.

OpenStudy (jamesj):

The important thing is (a/b)y' + y = 0 is a linear equation. Hence if y1 is a solution of the equation, then Ay1 is also a solution for any constant A.

OpenStudy (anonymous):

ahhh oke :)

OpenStudy (jamesj):

Hence for our purposes we could have supposed c = 0. Then we would have found a (i.e., one) solution \[ y = e^{-(b/a)x} \] Then the general solution of the equation is \[ y = Ae^{-(b/a)x} \] for an arbitrary constant A.

OpenStudy (jamesj):

It's very good practice though to not suppose c = 0, as a way to remind yourself there is a degree of freedom in the general solution. It is also essential when solving a number of inhomogeneous equations that the c stay in expressions.

myininaya (myininaya):

you mean x(t)?

OpenStudy (jamesj):

For example, in this equation is x(t) were not identically zero, then when working through the solution, you want to keep the c 'alive' in the algebra.

OpenStudy (jamesj):

but there are other sorts of inhomogeneous equations of course.

myininaya (myininaya):

oh i see what c you are talking about

OpenStudy (anonymous):

what about the particular solution... for example.. (a/b)y'+y = 1

myininaya (myininaya):

i love my way above

OpenStudy (jamesj):

In that case the particular solution is yp(x) = 1.

OpenStudy (anonymous):

could you explain why it is 1?

OpenStudy (jamesj):

Substitute and see: it works.

myininaya (myininaya):

get the coeificent of y' to be 1 and find integrating factor

myininaya (myininaya):

but do what james said

myininaya (myininaya):

i like to make things extra long

OpenStudy (jamesj):

In fact, for any constant K, the particular solution of (a/b)y' + y = K is yp(x) = K

OpenStudy (anonymous):

oke, thats clear, but what do you mean with substitute... what do i have to substitute?

myininaya (myininaya):

K=1

OpenStudy (anonymous):

owh wait... for an example.. if it was y'+y = x yp = ax+b right?

OpenStudy (anonymous):

then i have to get the first derivative

OpenStudy (anonymous):

and substitute it in y'+y

OpenStudy (jamesj):

I mean the function yp(x) = K. Substitute it into the equation and you see it indeed satisfies (a/b)y' + y = K

OpenStudy (jamesj):

For the method of solving (a/b)y' + y = x(t) in general, watch this: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-7-first-order-linear-with-constant-coefficients/

OpenStudy (anonymous):

oke jamesj.. im gonna watch the video :) thanks <3

OpenStudy (jamesj):

Actually, I now realize this is slightly too far along. Instead, start at lecture 3, here: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/

myininaya (myininaya):

\[y'+\frac{b}{a}y=x \frac{b}{a}\] \[vy'+v \frac{b}{a}y=v x \frac{b}{a}\] we want to choose v (v>0) such that \[v'=v \frac{b}{a} => \frac{dv}{dt}=v \frac{b}{a} => \frac{1}{v} dv=\frac{b}{a} dt \] now integrating both sides gives \[\ln(v)=\frac{b}{a} t +k \text{ let k=0} => \ln(v)=\frac{b}{a} t => v=e^{\frac{b}{a} t}\] we chose this v this way so that we could write vy'+v'y=(vy)' so we have \[(e^{\frac{ b}{a} t} y)'=e^{\frac{b}{a} t} x \frac{b}{a}\] \[(e^{\frac{ b}{ a} t} y)'=\frac{b}{a} x e^{\frac{b}{a} t}\] integrating both sides we get \[e^{\frac{b}{a}t} y=\int\limits_{}^{}\frac{b}{a}xe^{\frac{b}{a} t} d t+C\] \[y=\frac{b}{a e^{\frac{b}{a} t}} \int\limits_{}^{} x e^{\frac{b}{a} t} dt +C\] i made a formula for the general case :)

myininaya (myininaya):

this is what i meant to do way above but i wrote t instead of x

OpenStudy (anonymous):

ahh okee... thnks myininaya :)

myininaya (myininaya):

do you understand what i did? is this consider to advance?

myininaya (myininaya):

too*

OpenStudy (jamesj):

@myin: carefully in your last expression, as the exponential multiplies both the integral and the constant.

myininaya (myininaya):

oops

myininaya (myininaya):

james is write the last expression i should have wrote was \[y=\frac{b}{a e^{\frac{b}{a} t}} ( \int\limits\limits_{}^{} x e^{\frac{b}{a} t} dt +C )\]

OpenStudy (anonymous):

its a bit complex... but i think i got it :)

myininaya (myininaya):

above the constant wasn't being multiplied by b/a so it doesn't matter the way that i wrote it since b/a *C is still a constant

myininaya (myininaya):

but still the constant needs to be divided by that exponential guy

myininaya (myininaya):

its hard for some to use the product rule backwards and thats what i was trying to do above

myininaya (myininaya):

fg'+f'g = (fg)'

OpenStudy (anonymous):

yeahh.. indeed.. the product rule backwards is hard >.<

myininaya (myininaya):

so what if i asked you to write this as (something)' \[\frac{-1}{x^2} \cdot e^{-x^2}-2e^{-x^2}\] do you think you could do it?

OpenStudy (anonymous):

let me think

myininaya (myininaya):

maybe that was a hard one to start off with i should have chose an easier one like xy'+y

OpenStudy (anonymous):

oeehh this one is hard (x^-1*e^(x^2))'

myininaya (myininaya):

almost that x^2 should have negative in front it

myininaya (myininaya):

but excellent job

OpenStudy (anonymous):

yeaaaahh ^^ yeahh dammit i forgot that one -.-

myininaya (myininaya):

that is useful for solving things in this form y'+p(x)y=q(x) this is called a linear differential equation

myininaya (myininaya):

some might require you to use the formula but it is also good understand how the formula is derived

myininaya (myininaya):

i mean they have a formula for y'+p(x)y=q(x)

myininaya (myininaya):

i don't remember it i just derive it every single time

myininaya (myininaya):

using the long process above

OpenStudy (jamesj):

I forget the formula all the time, but one can rederive it in a minute. Anyway, it's all in the lecture 3 I linked to earlier: http://ocw.mit.edu/courses/mathematics/18-03-differential-equations-spring-2010/video-lectures/lecture-3-solving-first-order-linear-odes/

myininaya (myininaya):

sometimes you can manipulate a nonlinear differential equation so you can use the process i have above

myininaya (myininaya):

i like doing that

myininaya (myininaya):

i have only done it once though lol